### v3

# RIKEN-iTHEMS-Report-25, KYUSHU-HET-308

Abstract
The ’t Hooft partition function $\mathcal{Z}_{\text{tH}}[E;B]$ of an $SU(N)$ gauge theory with the $\mathbb{Z}_{N}$ 1-form symmetry is defined as the Fourier transform of the partition function $\mathcal{Z}[B]$ with respect to the spatial-temporal components of the ’t Hooft flux $B$. Its large volume behavior detects the quantum phase of the system. When the integrand of the functional integral is real-positive, the latter partition function $\mathcal{Z}[B]$ can be numerically computed by a Monte Carlo simulation of the $SU(N)/\mathbb{Z}_{N}$ gauge theory, just by counting the number of configurations of a specific ’t Hooft flux $B$. We carry out this program for the $SU(2)$ pure Yang–Mills theory with the vanishing $\theta$-angle by employing a newly-developed hybrid Monte Carlo (HMC) algorithm (the halfway HMC) for the $SU(N)/\mathbb{Z}_{N}$ gauge theory. The numerical result clearly shows that all non-electric fluxes are “light” as expected in the ordinary confining phase with the monopole condensate. Invoking the Witten effect on $\mathcal{Z}_{\text{tH}}[E;B]$, this also indicates the oblique confinement at $\theta=2\pi$ with the dyon condensate.
[…]
B01,B03,B31
1 Introduction
For an $SU(N)$ gauge theory with the $\mathbb{Z}_{N}$ 1-form symmetry Gaiotto:2014kfa , such as the pure Yang–Mills theory or the $\mathcal{N}=1$ and $\mathcal{N}=1^{*}$ supersymmetric Yang–Mills theories Donagi:1995cf , one can introduce the ’t Hooft flux $B:=(B_{12},B_{13},B_{14},B_{23},B_{24},B_{34})\in\mathbb{Z}_{N}$ by the twisted boundary conditions on the 4 torus $T^{4}$ tHooft:1979rtg ; tHooft:1981sps . The ’t Hooft partition function $\mathcal{Z}_{\text{tH}}[E;B]$ is then defined as the Fourier transform of the partition function $\mathcal{Z}[B]$ with respect to the spatial-temporal components of the ’t Hooft flux $B$ tHooft:1979rtg ; tHooft:1981sps :
$\mathcal{Z}_{\text{tH}}[E_{1},E_{2},E_{3};B_{12},B_{23},B_{31}]:=\frac{1}{N^{3}}\sum_{B_{14},B_{24},B_{34}=0}^{N-1}\exp\left(\frac{2\pi i}{N}\sum_{i=1}^{3}E_{i}B_{i4}\right)\mathcal{Z}[B].$ (1.1)
On the left-hand side, $E_{i}$ and $B_{ij}$ are referred to as the electric and magnetic fluxes, respectively. The large volume behavior of $\mathcal{Z}_{\text{tH}}[E;B]$ detects the quantum phase (i.e., confinement, Higgs, or Coulomb) of the system tHooft:1979rtg ; tHooft:1981sps . See also Ref. Tomboulis:1985ah . Recently, consideration of the ’t Hooft partition function $\mathcal{Z}_{\text{tH}}[E;B]$ has been revived Kitano:2017jng ; Tanizaki:2022ngt ; Nguyen:2023fun , largely motivated by the perspective of the generalized symmetries Gaiotto:2014kfa , in particular in connection with the study in Ref. Gaiotto:2017yup ; see also Refs. Yamazaki:2017ulc ; Hayashi:2023wwi ; Hayashi:2024qkm ; Hayashi:2024yjc ; Hayashi:2024psa for recent related studies.
Now, when the integrand of the functional integral is real-positive as the case in the pure Yang–Mills theory with the vanishing $\theta$-angle, the partition function $\mathcal{Z}[B]$ on the right-hand side of Eq. (1.1), or more precisely the ratio $\mathcal{Z}[B]/\mathcal{Z}[0]$, may be numerically computed by a Monte Carlo simulation Kovacs:2000sy ; deForcrand:2001nd . Traditionally, this ratio is computed by “reweighting” the difference of lattice actions with and without $B$ Kovacs:2000sy ; deForcrand:2001nd . In the present paper, we employ a Monte Carlo simulation of the $SU(N)/\mathbb{Z}_{N}$ Yang–Mills theory on the basis of a recently developed hybrid Monte Carlo (HMC) algorithm (the halfway HMC) Abe:2024fpt in which the ’t Hooft flux $B$ (the total flux of the $\mathbb{Z}_{N}$ 2-form flat gauge field) is explicitly treated as a dynamical variables. [^1] For a more “traditional” approach to the $SU(N)/\mathbb{Z}_{N}$ Yang–Mills theory using the plaquette action in the adjoint representation, see Refs. Halliday:1981te ; Creutz:1982ga . See also Refs. Edwards:1998dj ; deForcrand:2002vs ; Halliday:1981tm . In this way, each configuration generated in the Monte Carlo simulation possesses a various but definite value of the ’t Hooft flux $B$. Then, just by counting the number of configurations of a specific $B$, we can obtain the partition function $\mathcal{Z}[B]$. [^2] We expect that our computational method is free from the overlap problem that requires an elaborate computational trick in the reweighting approach Kovacs:2000sy ; deForcrand:2001nd . The ’t Hooft partition function $\mathcal{Z}_{\text{tH}}[E;B]$ is then given by Eq. (1.1). Explicitly, we carry out this program for the $SU(2)$ pure Yang–Mills theory with a vanishing $\theta$-angle. [^3] In this paper, we study the partition function in the presence of the ’t Hooft flux, not the ’t Hooft line tHooft:1977nqb . The former is the total flux of the $\mathbb{Z}_{N}$ 2-form flat gauge field such that $\mathrm{d}B=0\bmod N$ (i.e., elements of $H^{2}(T^{4},\mathbb{Z}_{N})$), while for the latter $\mathrm{d}B$ is given by the Poincaré dual of the line Aharony:2013hda . Computationally, the study of the latter Hoelbling:1998gv ; Hoelbling:1999gr ; Hoelbling:2000su ; deForcrand:2000fi ; DelDebbio:2000cb is more demanding. Our numerical result below clearly shows that all nonelectric fluxes are “light” as expected in the ordinary confining phase with the monopole condensate tHooft:1979rtg ; tHooft:1981sps . Although the study of the ’t Hooft partition function has a long history, to our knowledge, this is the first attempt to measure the ’t Hooft partition function with all possible combinations of the ’t Hooft flux by a lattice Monte Carlo simulation.
As in our numerical calculation, when all cycles of the 4 torus $T^{4}$ possess an equal radius $L$, the partition function $\mathcal{Z}[B]$ enjoys the Euclidean $90^{\circ}$ rotational invariance (see Appendix A) and, as the consequence of this, $\mathcal{Z}_{\text{tH}}[E;B]$ obeys the duality equation tHooft:1979rtg ; tHooft:1981sps ,
$\mathcal{Z}_{\text{tH}}[E_{1},E_{2},E_{3};B_{12},B_{23},B_{31}]$
$=\frac{1}{N^{2}}\sum_{B_{23}^{\prime},B_{31}^{\prime},E_{1}^{\prime},E_{2}^{\prime}=0}^{N-1}\exp\left[\frac{2\pi i}{N}\left(E_{1}B_{23}^{\prime}+E_{2}B_{31}^{\prime}-B_{23}E_{1}^{\prime}-B_{31}E_{2}^{\prime}\right)\right]$
×$\qquad\qquad\qquad\qquad\qquad{}\times{}\mathcal{Z}_{\text{tH}}[E_{1}^{\prime},E_{2}^{\prime},E_{3};B_{12},B_{23}^{\prime},B_{31}^{\prime}].$ (1.2)
We observe that to fairly good numerical accuracy our numerical result for $\mathcal{Z}_{\text{tH}}[E;B]$ fulfills this equation, providing a consistency check of the computation. We may even take the average of $\mathcal{Z}[B]$ over Euclidean $90^{\circ}$ rotations so that $\mathcal{Z}_{\text{tH}}[E;B]$ automatically fulfills this duality equation within the numerical error (see below).
2 Direct computation of the ’t Hooft partition function
Our lattice action on a periodic lattice of size $L$, $\Gamma:=(\mathbb{Z}/L\mathbb{Z})^{4}$, for the $SU(N)/\mathbb{Z}_{N}$ theory is given by Mack:1978kr ; Ukawa:1979yv ; Srednicki:1980gb ; Seiler:1982pw (see also Ref. Kapustin:2014gua ):
[…]
where $\beta$ is the bare coupling and the plaquette variables $P(x,\mu,\nu)$ are given by
$P(x,\mu,\nu):=U(x,\mu)U(x+\hat{\mu},\nu)U(x+\hat{\nu},\mu)^{\dagger}U(x,\nu)^{\dagger}$ (2.2)
from $SU(N)$ link variables. The integer field $B_{\mu\nu}(x)$ in Eq. (2.1) is given by the ’t Hooft flux by
$B_{\mu\nu}(x)=\begin{cases}B_{\mu\nu}&\text{for $x_{\mu}=L-1$ and~{}$x_{\nu}=L-1$},\\ 0&\text{otherwise}\\ \end{cases}$ (2.3)
with $B_{\mu\nu}=\{0,1,\dotsc,N-1\}\bmod N$ (we set $B_{\nu\mu}=-B_{\mu\nu}$). For the Boltzmann weight $e^{-S}$, we generated configurations of $(U,B)$ for $N=2$ by employing the halfway HMC. We refer the reader to Ref. Abe:2024fpt for details of our numerical simulation. [^4] Our numerical codes can be found in https://github.com/o-morikawa/Gaugefields.jl, which is based on Gaugefields.jl in the JuliaQCD package Nagai:2024yaf . The partition function $\mathcal{Z}[B]$ with a particular ’t Hooft flux, say, $(B_{12},B_{13},B_{14},B_{23},B_{24},B_{34})=(0,0,0,1,1,1)$ can be obtained as the expectation value of the operator,
$\mathcal{O}_{(0,0,0,1,1,1)}:=\frac{1}{2^{6}}\delta_{B,(0,0,0,1,1,1)}.$ (2.4)
The expectation value, however, can be computed by simply counting the number of configurations with the flux $(0,0,0,1,1,1)$ and dividing it by the total number of configurations. [^5] We store gauge field configurations generated by the halfway HMC Abe:2024fpt under filenames such as U_beta2.6_L20_F111001_7020.txt, where the number 111001 represents the value of the flux $(B_{12},B_{13},B_{14},B_{23},B_{24},B_{34})$ of that configuration.
We may further consider the average of $\mathbb{Z}[B]$ over the Euclidean $90^{\circ}$ orthogonal rotations. See Appendix A for a list of the irreducible representations. For instance, corresponding to the dimension 4 irreducible representation in Eq. (A.8), Eq. (2.4) may be replaced by
$\bar{\mathcal{O}}_{(0,0,0,1,1,1)}:=\frac{1}{2^{6}}\frac{1}{4}\left[\delta_{B,(0,0,0,1,1,1)}+\delta_{B,(0,1,1,0,0,1)}+\delta_{B,(1,0,1,0,1,0)}+\delta_{B,(1,1,0,1,0,0)}\right].$ (2.5)
We will see that this prescription reduces the statistical error considerably.
In what follows, we show the results using 2590 configurations for $\beta=2.6$ and $L=20$. No attempt to find the continuum limit is made because the behavior is almost the same for all lattice parameters considered in Ref. Abe:2024fpt .
First, in Fig. 1, we plot the ’t Hooft partition function $\mathcal{Z}_{\text{tH}}[E;B]$ for all possible combinations of ’t Hooft fluxes, $E_{i}$ and $B_{ij}$. The statistical errors are estimated by the jackknife method. [^6] We found that statistical errors are almost saturated at the bin size $\sim 1$. We think that this is a consequence of a shortness of the autocorrelation length in the present HMC algorithm Abe:2024fpt . In this figure, no average over Euclidean $90^{\circ}$ rotations is taken. The filled symbols represent $\mathcal{Z}_{\text{tH}}[E;B]$ computed from $\mathcal{Z}[B]$ by Eq. (1.1) and the unfilled symbols represent the right-hand side of the duality equation (1.2). We thus observe that the duality equation (1.2) holds to a fairly good accuracy.
Figure 1: The ’t Hooft partition function $\mathcal{Z}_{\text{tH}}[E;B]$ for all possible combinations of the ’t Hooft fluxes, $E_{i}$ and $B_{ij}$. $\beta=2.6$ and $L=20$. The statistical errors are estimated by the jackknife method. No average over Euclidean $90^{\circ}$ rotations is taken. The filled symbols represent $\mathcal{Z}_{\text{tH}}[E;B]$ computed from $\mathcal{Z}[B]$ by Eq. (1.1) and the unfilled symbols represent the right-hand side of the duality equation (1.2). We observe that the duality equation holds to a fairly good accuracy.
In Ref. Nguyen:2023fun , the authors showed that the ’t Hooft partition function $\mathcal{Z}_{\text{tH}}[E;B]$ is real semi-positive as far as the reflection positivity holds. The plots in Fig. 1 are also consistent with this property within the errors.
From Fig. 1, we immediately observe that the ’t Hooft partition functions $\mathcal{Z}_{\text{tH}}[E;B]$ are clearly classified into two classes depending on the value of the flux; one gives $\mathcal{Z}_{\text{tH}}[E;B]/\mathcal{Z}_{\text{tH}}[E=0;B=0]\sim 1$ and another gives $\mathcal{Z}_{\text{tH}}[E;B]/\mathcal{Z}_{\text{tH}}[E=0;B=0]\sim 0$. The former class of fluxes is called “light” and the latter class is called “heavy” tHooft:1979rtg ; tHooft:1981sps . From the duality equation (1.2), it can be argued that, when $E_{3}$ and $B_{12}$ are fixed, there are only 00 or $N^{2}=4$ combinations of light fluxes among totally $N^{4}=16$ combinations of fluxes tHooft:1979rtg ; tHooft:1981sps . See also Ref. Nguyen:2023fun . We clearly see that this assertion holds in Fig. 1. We also observe that all fluxes with $E_{i}\neq 0$ are heavy. In other words, no light fluxes possess electric flux, $E_{i}=0$. This is expected in the ordinary confining phase in which the magnetic monopole condensates tHooft:1979rtg ; tHooft:1981sps . [^7] The inverse Fourier transform of $\mathcal{Z}_{\text{tH}}[E;B]$ with this behavior in the large volume limit gives $\mathcal{Z}[B]/\mathcal{Z}[B=0]\sim 1$ for any $B$. In modern language Gaiotto:2014kfa ; Nguyen:2023fun , this implies that one can define the $\mathbb{Z}_{N}$ 1-form symmetry in the low-energy theory, whose symmetry operator is spanned by a 2-surface given by the Poincaré dual of a given flux $B$. When $B_{34}=1$ and other components vanish, the symmetry operator is spanned by a 2-surface in the $12$-direction. Then a move of a fundamental Wilson line extending in the $4$-direction along the cycle in the $3$-direction produces the factor $e^{\pm 2\pi i/N}$. This factor forbids the expectation value of the Wilson line (or the Polyakov line) and implies that the $\mathbb{Z}_{N}$ 1-form symmetry is not spontaneously broken; this is a characterization of the ordinary confining phase.
Figure 2 is the same as Fig. 1, but the average over the Euclidean $90^{\circ}$ rotations is taken in $\mathcal{Z}[B]$. The statistical errors become smaller as anticipated; the duality equation is automatically fulfilled and the separation into light fluxes and heavy fluxes becomes clearer.
Figure 2: The ’t Hooft partition function $\mathcal{Z}_{\text{tH}}[E;B]$ for all possible combinations of ’t Hooft fluxes, $E_{i}$ and $B_{ij}$. $\beta=2.6$ and $L=20$. The errors are estimated by the jackknife method. The average of $\mathcal{Z}[B]$ over Euclidean $90^{\circ}$ rotations is made.
Our result thus illustrates that a direct numerical computation of the ’t Hooft partition function can become a useful approach to study the quantum phase of $SU(N)$ gauge theories with the $\mathbb{Z}_{N}$ 1-form symmetry, at least as far as the integrand of the functional integral is real-positive; when the complex phase of the integrand is small, it could be treated by reweighting.
The behavior $\mathcal{Z}_{\text{tH}}[E;B]/\mathcal{Z}_{\text{tH}}[E=0;B=0]\to 1$ or $\mathcal{Z}_{\text{tH}}[E;B]/\mathcal{Z}_{\text{tH}}[E=0;B=0]\to 0$ is an asymptotic one in the large volume limit tHooft:1979rtg ; tHooft:1981sps and the way of approaching it provides useful information on the (dual) string tension. Unfortunately, we found that it is difficult at the moment to determine the sting tension in our simulation with the present lattice parameters and statistics.
We may generalize the partition function as $\mathcal{Z}_{\theta}[B]$ by multiplying the $\theta$-term $e^{-i\theta Q}$ to the Boltzmann weight. In the pure $SU(N)$ Yang–Mills theory, $Q\in\mathbb{Z}$, but in the presence of the ’t Hooft flux, $Q$ becomes fractional as tHooft:1981nnx ; vanBaal:1982ag :
$Q=-\frac{1}{N}\frac{\varepsilon_{\mu\nu\rho\sigma}B_{\mu\nu}B_{\rho\sigma}}{8}+\mathbb{Z}.$ (2.6)
It is possible to construct a geometric definition of $Q$ on the lattice that possesses this property Abe:2023ncy by requiring the $\mathbb{Z}_{N}$ 1-form gauge symmetry in the construction in Ref. Luscher:1981zq . Noting Eq. (2.6), under the shift $\theta\to\theta+2\pi$, one finds that the corresponding ’t Hooft partition function $\mathcal{Z}_{\text{tH},\theta}[E;B]$ (1.1) behaves as
$\mathcal{Z}_{\text{tH},\theta+2\pi}[E_{1},E_{2},E_{3};B_{12},B_{23},B_{31}]=\mathcal{Z}_{\text{tH},\theta}[E_{1}+B_{23},E_{2}+B_{31},E_{3}+B_{12};B_{12},B_{23},B_{31}].$ (2.7)
This is the Witten effect Witten:1979ey on the ’t Hooft partition function. Invoking this relation with $\theta=0$, from Figs. 1 and 2, we infer that fluxes with $E_{1}+B_{23}=E_{2}+B_{31}=E_{3}+B_{12}=0\bmod 2$ are light for $\theta=2\pi$; see Fig. 3 for the ’t Hooft partition function $\mathcal{Z}_{\text{tH},\theta=2\pi}[E;B]$. This pattern of light fluxes is an indication of the oblique confinement, in which the dyons condensate tHooft:1979rtg ; tHooft:1981bkw ; tHooft:1981sps .
Figure 3: The ’t Hooft partition function with $\theta=2\pi$, $\mathcal{Z}_{\text{tH},\theta=2\pi}[E;B]$ for all possible combinations of ’t Hooft fluxes, $E_{i}$ and $B_{ij}$. $\beta=2.6$ and $L=20$. This plot was obtained by simply applying the Witten effect (2.7) with $\theta=0$ to data in Fig. 2.
3 Conclusion
The pattern exhibited by the ’t Hooft partition function $\mathcal{Z}_{\text{tH}}[E;B]$ as the function of the flux, such as that in Figs. 2 and 3, clearly indicates the quantum phase of the system tHooft:1979rtg ; tHooft:1981sps . In this paper, by employing the halfway HMC Abe:2024fpt for the simplest example, the $SU(2)$ pure Yang–Mills theory, we illustrated that a direct Monte Carlo calculation of the pattern is feasible, at least as long as the integrand of the functional integral is real-positive; when the complex phase of the integrand is small, it could be treated by reweighting. Our methodology itself is quite general so we hope to carry out analyses in more intriguing situations: It should be interesting to study finite temperature cases (such as in Ref. deForcrand:2001nd ) to observe the confining/deconfining phase transition. The inclusion of matter fields such as the adjoint Higgs and adjoint fermions will give a more intricate pattern of the ’t Hooft partition function. We hope to tackle these problems in the near future.
Acknowledgments
We would like to thank Motokazu Abe for discussions spanning a long period of time, which finally led to this work. We also thank Philippe de Forcrand, Yui Hayashi, Yuki Nagai, Soma Onoda, Yuya Tanizaki, Akio Tomiya, and Hiromasa Watanabe for helpful discussions. We appreciate the opportunity of discussion during the YITP–RIKEN iTHEMS conference “Generalized symmetries in QFT 2024” (YITP-W-24-15) in the execution of this work. The numerical computations in this paper were carried out on Genkai, a supercomputer system of the Research Institute for Information Technology (RIIT), Kyushu University. O.M. acknowledges the RIKEN Special Postdoctoral Researcher Program. The work of H.S. was partially supported by a Japan Society for the Promotion of Science (JSPS) Grant-in-Aid for Scientific Research, JP23K03418.
Appendix A Euclidean $90^{\circ}$ discrete rotations on the $\mathbb{Z}_{2}$ flux in 4D and 3D
On the components of a second rank antisymmetric tensor,
$B_{1}:=B_{12},\quad B_{2}:=B_{13},\quad B_{3}:=B_{14},\quad B_{4}:=B_{23},\quad B_{5}:=B_{24},\quad B_{6}:=B_{34},$ (A.1)
Euclidean $90^{\circ}$ rotations on the $\mu\nu$-plane are represented as $B_{a}\to\sum_{b=1}^{6}(\Lambda_{\mu\nu})_{ab}B_{b}$, where the representation matrices are given by,
$\Lambda_{12}=\begin{pmatrix}1&0&0&0&0&0\\ 0&0&0&-1&0&0\\ 0&0&0&0&-1&0\\ 0&1&0&0&0&0\\ 0&0&1&0&0&0\\ 0&0&0&0&0&1\\ \end{pmatrix},\qquad\Lambda_{13}=\begin{pmatrix}0&0&0&1&0&0\\ 0&1&0&0&0&0\\ 0&0&0&0&0&-1\\ -1&0&0&0&0&0\\ 0&0&0&0&1&0\\ 0&0&1&0&0&0\\ \end{pmatrix},$
$\Lambda_{14}=\begin{pmatrix}0&0&0&0&1&0\\ 0&0&0&0&0&1\\ 0&0&1&0&0&0\\ 0&0&0&1&0&0\\ -1&0&0&0&0&0\\ 0&-1&0&0&0&0\\ \end{pmatrix},\qquad\Lambda_{23}=\begin{pmatrix}0&-1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 0&0&0&0&0&-1\\ 0&0&0&0&1&0\\ \end{pmatrix},$
$\Lambda_{24}=\begin{pmatrix}0&0&-1&0&0&0\\ 0&1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&0&0&1\\ 0&0&0&0&1&0\\ 0&0&0&-1&0&0\\ \end{pmatrix},\qquad\Lambda_{34}=\begin{pmatrix}1&0&0&0&0&0\\ 0&0&-1&0&0&0\\ 0&1&0&0&0&0\\ 0&0&0&0&-1&0\\ 0&0&0&1&0&0\\ 0&0&0&0&0&1\\ \end{pmatrix}.$ (A.2)
Some configurations of the ’t Hooft flux are thus related by these transformations. For $N=2$, for which $B_{a}=\{0,1\}\bmod 2$, the configurations are classified into the following 11 irreducible representations, Eqs. (A.3)–(A.9) and combinations obtained by the interchange $0\leftrightarrow 1$ in Eqs. (A.3)–(A.6):
$(B_{1},B_{2},B_{3},B_{4},B_{5},B_{6})=(0,0,0,0,0,0),$ (A.3)
$(0,0,0,0,0,1),\quad(0,0,0,0,1,0),\quad(0,0,1,0,0,0),$ (A.4a)
$(0,0,0,1,0,0),\quad(0,1,0,0,0,0),\quad(1,0,0,0,0,0),$ (A.4b)
$(0,0,0,0,1,1),\quad(0,0,1,0,0,1),\quad(0,0,1,0,1,0),$ (A.5a)
$(0,1,0,1,0,0),\quad(1,0,0,1,0,0),\quad(1,1,0,0,0,0),$ (A.5b)
$(0,0,0,1,0,1),\quad(0,0,0,1,1,0),\quad(0,1,0,0,0,1),$
$(0,1,1,0,0,0),\quad(1,0,0,0,1,0),\quad(1,0,1,0,0,0),$ (A.5c)
$(0,0,1,1,0,0),\quad(0,1,0,0,1,0),\quad(1,0,0,0,0,1),$ (A.6)
$(0,0,1,0,1,1),$ (A.7a)
$(0,1,0,1,0,1),\quad(1,0,0,1,1,0),\quad(1,1,1,0,0,0),$ (A.7b)
$(1,1,0,1,0,0),$ (A.8a)
$(0,0,0,1,1,1),\quad(0,1,1,0,0,1),\quad(1,0,1,0,1,0),\quad$ (A.8b)
$(0,0,1,1,0,1),\quad(0,0,1,1,1,0),\quad(0,1,0,0,1,1),$
$(0,1,1,0,1,0),\quad(1,0,0,0,1,1),\quad(1,0,1,0,0,1),$ (A.9a)
$(0,1,0,1,1,0),\quad(0,1,1,1,0,0),\quad(1,0,0,1,0,1),$
$(1,0,1,1,0,0),\quad(1,1,0,0,0,1),\quad(1,1,0,0,1,0).$ (A.9b)
We may thus take the average of the partition function $\mathcal{Z}[B]$ over elements within each of these irreducible representations. This average is taken in Figs. 2 and 3.
For finite temperatures, $\mathcal{Z}[B]$ is invariant only under 3D spatial $90^{\circ}$ rotations generated by $\Lambda_{12}$, $\Lambda_{13}$, and $\Lambda_{23}$. Then each irreducible representation is decomposed into smaller multiplets labeled by a roman index in the equation number, such as “a” in (A.5a). We may then take the average of $\mathcal{Z}[B]$ over elements within these smaller multiplets.